Observers are: V. Testa, S. Herbert-Fort, D. Gonzales Huerta

Mostly sunny with few thin clouds in the afternoon

00:35 - 25 bias

003632-005038

01:00 - Sunset

Still ongoing work in the dome. Flat fields delayed.

01:21 - Move to the FF spot.

Try B, for which no flats have yet been taken

01:25 - Sky Flat Test B: 012719
Counts/sec : 3300

01:28 - SkyFlatB_5 (exp.time = 15.0s)

012833-013407

022833: counts=31954
012958: " =23850
013127: " =17963
013238: " =14371
013407: " =10827

01:35 - SkyFlatBrot180_5 (exp.time = 30.0s)

013724-

013724: counts=11753
013902: " =8784
014029: " =6821
014200: " =5217
014333: " =3987

01:45 - Vfastextra on the flat field spot

Image Analysis: 014705

Median external diameter: 55.91 px
WARNING: Z11 TOO LARGE
Truncating value to 1000
% Compiled module: WRITEOUTFILE.
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-979 -352 -435 -2649 1721 1000
-----------------------------------------
Sent to TCS. Retry.

01:49 - Image Analysis: 015049
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 56.54 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-864 -455 -9 323 -575 92
-----------------------------------------
Sent. Moved to SA110_L1 field.

01:53 - Move to SA110_L1 and do IA again
Image Analysis: 015654
Median external diameter: 68.72 px
WARNING: Z11 TOO LARGE
Truncating value to 1000
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
1341 346 -966 2263 -806 1000
-----------------------------------------
Sent to TCS. Redone, probably seeing not too exciting

01:58 - Image Analysis: -----
Telescope stopped tracking.
OB stopped.

02:02 - Tried sending the OB again. Interface stuck, not playing the OB....
David Gonzales is trying to restore operations

02:04 - previously sent OB started. Stopped because telescope not tracking

02:09 - Telescope seems to track again.
Preset to SA110_L1 again
Vfastextra

02:10 - Lost tracking again. OB stopped.
Telescope has balancing problems.
Intervention made on the software and hardware
More weights added to balance the telescope.

03:50 - Ready to restart observations. Testing the telescope.
03:56 - Tracking ok. Raise the mirror.
Problems reported, called Tucson to fix
the mirror.

04:45 - Mirror seems ok.

04:46 - Move to AndXIV_V3 and Vfastextra

Image Analysis: 044918
----------------------------------------------------
No GOOD pupils found... check input image!
----------------------------------------------------
Try resetting the coefficients and try again.

04:52 - Image Analysis: 045322
----------------------------------------------------
No GOOD pupils found... check input image!
----------------------------------------------------
Try another one after making a collimation

04:55 - Image Analysis: 045511
----------------------------------------------------
No GOOD pupils found... check input image!
----------------------------------------------------
Call John Hill for advice on how to deal with this challenging run of IA.

He suggests, after clearing coefficients:

Z5 +3000
Z7 -2000
Z11 +200

05:00 - LBC crashed meanwhile

05:02 - Try again an
Image Analysis: 050242
WARNING: Z4 TOO LARGE, will consider defocus only
all other aberration coefficients set zero.
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-2616 0 0 0 0 0
-----------------------------------------
Sent to TCS and given another Z7 -1000 and tried again

05:05 - Image Analysis: 050553
----------------------------------------------------
No GOOD pupils found... check input image!
----------------------------------------------------
Z11 +200

05:08 - Image Analysis: 050843
----------------------------------------------------
No GOOD pupils found... check input image!
----------------------------------------------------

05:59 - Try with a longer exposure (32.0s)
Image Analysis: 051105
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 60.35 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-174 1190 2621 2897 906 666
-----------------------------------------
Sent to TCS and retry

05:12 - Image Analysis:051408
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 56.86 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-806 2179 294 -1649 -624 -132
-----------------------------------------
Sent and repeated.

05:16 - Image Analysis: 051700
Median external diameter: 27.53 px
WARNING: Z4 TOO LARGE, will consider defocus only
all other aberration coefficients set zero.
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-6121 0 0 0 0 0
-----------------------------------------
Not sent, move to another field.

Go back to the ordinary 16.0s exposure time.

05:20 - Image Analysis: 052037
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 59.64 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-303 -1699 -811 89 253 171
-----------------------------------------
Sent to TCS and retry

05:23 - Image Analysis: 052314
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 60.20 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-201 -691 580 -95 -15 72
-----------------------------------------
Sent to TCS and retry

05:25 - Image Analysis: 052503
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 60.34 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-175 268 56 -208 -32 5
-----------------------------------------
Sent to TCS and move to scientific field

05:26 - AndXIV_V3

052826-054025

052826: seeing = 1.2"
053432: " = 1.2"
054025: " = 1.3"

05:46 - AndXIV_B3

054718-055932

054718: seeing = 1.3"
055322: " = 1.3"
055932: " = 1.4"

06:05 - Vfastextra

Image Analysis: 060634
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 58.34 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-538 493 273 333 339 96
-----------------------------------------
Sent to TCS

06:08 - AndXIV_V4

060915-********

060915: seeing = stars appear double, as happened with 2006 jumps.
061522: " = aborted, reset the encoders and redo the OB

06:25 - Reset the encoders, tried loading Vfastextra cmu did not accept
because of disk full. Moved some old OBs to another filesystem
and retried.

06:28 - Image Analysis: OB does not start, interface stuck
started at its best convenience: 063749
seems like a long timeout...
exposed while telescope moving, no pupils obviusly found.

06:39 - Try again
Image Analysis: 063957
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 56.31 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-906 -552 -807 16 81 -133
-----------------------------------------
Sent to TCS, then again the science OB

06:41 - AndXIV_V4 again
Telescope stopped again in elevation

07:16 - We have the telescope elevation back in work

Sent a preset to AndXIV_V4 and done Image Analysis.

07:17 - Image Analysis: 071847

Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 59.96 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-244 -417 -487 166 -2 -12
-----------------------------------------

Sent to TCS and start the OB

07:22 - AndXIV_V4, try again

072420-073652

072420: seeing = 1.5"
073035: seeing = 1.2" (but made another jump)
073652: seeing = 1.3"

07:44 - AndXIV_B4

074503-075721

074503: seeing = 1.55"
075109: " = 1.6"
075721: " =

08:05 - Vfastextra: 080633

Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 62.02 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
128 281 -65 -9 24 -133
-----------------------------------------
Sent to TCS and go to target

08:08 - AndXIV_V5

080857-082103

080857: seeing = 1.3"
081450: " = 1.3"
082103: " = 1.5"

08:28 - AndXIV_B5

082930-084127

082930: seeing = 1.4"
083522: " = 1.6"
084127: " = 1.5"

08:47 - Vfastextra

Image Analysis: 084841
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 64.26 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
533 115 471 -123 29 99
-----------------------------------------
Sent to TCS. Continue on target

08:50 - AndXIV_V6

085124-090313

085124: seeing = 1.05"
085720: " = 1.0"
090313: " = 1.0"

09:11 - AndXIV_B6

091154-092342

091154: seeing = 1.0"
091747: seeing = 1.0"
092342: seeing = 0.9"

09:29 - Image Analysis: 093039
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 61.14 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-31 573 141 126 -24 154
-----------------------------------------
Sent to TCS. GO for the last 2 V,B OBs

09:32 - AndXIV_V7

093314-094450

093314: seeing = 1.0"
093859: " = 0.8"
094450: " = 0.9"

09:50 - AndXIV_B7

095051-100225

095051: seeing = 1.0"
095640: " = 1.0"
100225: " =

10:09 - Move to Clementini S2B

Image Analysis: 101057
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 59.48 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-332 555 192 -49 -50 -28
-----------------------------------------
Sent to TCS. Move to S2B1

10:12 - M31_S2B_1

101302-102430

101302: seeing = 1.2"
101851: seeing = 1.4"
102430: seeing = 1.3"

10:31 - Move to Stanek target

Image Analysis: 103356
WARNING: 1 star only used for computation!!
Computing aberrations using 1 pupils
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 61.13 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-32 -122 229 -24 167 244
-----------------------------------------

10:35 - StanekM33 : 1 (300x5)

103851-

103851: seeing = 1.3"
104437: " = 1.2"
105019: " = 1.25"
105558: " = 1.2"
110137: " = 1.3"

10:07 - Move to Kochanek M81

Image Analysis with OB Bfastextra special by CSK:
(Airmass=1.68) 111125
Focus Offset = -0.80000001 Filter = V-BESSEL
Median external diameter: 59.04 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
-411 402 165 -328 -478 -304
-----------------------------------------
Sent to TCS.

11:14 - KochanekM81BV

111531-112706

111531: seeing = 1.7"
111710:
111849:
112029:
112209:
112348:
112527:
112706:

11:30 - Played again:

113036-

113036:
113215:
113355:
113534:
113713:
113851:
114030:
114209: seeing = 1.6"

11:44 - Move to Nonino MS0451_C_2
Image Analysis: 114957

Focus Offset = -0.80000001 Filter = V-BESSEL(?)
Median external diameter: 64.32 px
-----------------------------------------
Computed Aberrations (nm)
Defoc AstX AstY ComaX ComaY Sph
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
-----------------------------------------
545 -526 -1143 630 8 111
-----------------------------------------
Sent to TCS. Start OB.

11:51 - MS0451_C_2
115304-121257

115304: seeing = 1.3"
115754: " =
120253: " =
120752: " =
121257: " = 1.3"

That's it for tonight. Let's go for standards....

SA95 & SA98

05:17 - Move to SA95_105 (UsUBVgr) 4 positions

121950-123829

Too bright for the other standard field

05:41 - Move to flat field spot, 1h W of the meridian

05:43 - SkyFlatTestVrot180 (exp.time = 1.0s)
124622: counts = 4000

05:47 - SkyFlatVrot180_5 (exp.time = 6.0s)

124747: counts = 36191
124833: counts = 43572
124925: counts = 53995
125010: counts = 64718 (!)
125055: counts = 64787 (!)

05:51 - SkyFlatTestBrot180
125200: counts = 13000

05:53 - SkyFlatBrot180_5 (exp.time = 2.0s)

125313: counts = 35231.
125354: counts = 42445.
125441: counts = 52443.
125522: counts = 62655. (!)
(rescaled exp.time to 1.0s)
125604: counts = 37423.

05:56 - SkyFlatUspecrot180_5 (exp.time = 1.0s)

125746: counts = 16114
125827: counts = 19188
125910: counts = 23203
125950: counts = 27433
130030: counts = 32559

13:00 - SkyFlatUspec_5 (exp.time = 1.0s)

130203: counts = 47603
130243: counts = 55889
130323: counts = 64584 (!)
130403: counts = 64788 (!)
130443: counts = 64857 (!)

13:07 - 25 bias

Written by V. Testa.

-- JohnHill - 11 Oct 2007
Topic revision: r1 - 11 Oct 2007, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback